REU: SCHEDULE FOR WEEKS SEVEN AND EIGHT

TALKS ARE IN ECKHART 206 (unless otherwise posted) (May’s morning talks in R276)

Ex–YSP participants are expected to arrange mentoring sessions with graduate students and problem–solving group sessions between 9:30 and 4:00. This is time free for you to work on problems.

Date  SESAME  Babai  Bader  Kerr  May
Mon, August 1 9–12, 1–4 3:00 – 4:45  5:00 – 6:00 11:00-12:00
Tue, August 2 9–12, 1–4  4:30 – 6:00
Wed, August 3 9–12, 1–4 3:00 – 4:45  5:00 – 6:00 11:00-12:00
Thu, August 4 9–12, 1–4  4:30 – 6:00
Fri, August 5 9–12, 1–4 3:00 – 4:45 5:00 – 6:00 11:00-12:00
Mon, August 8 9–12, 1–4 3:00 – 4:45  5:00 – 6:00 11:00-12:00
Tue, August 9 9–12, 1–4  4:30–6:00
Wed, August 10 9–12, 1–4 3:00 – 4:45  5:00 – 6:00 11:00-12:00
Thu, August 11 9–12, 1–4  4:30–6:00
Fri, August 12 9–12, 1–4 3:00 – 4:45  5:00 – 6:00 11:00-12:00

Here is the schedule in summary form:

SEVENTH WEEK:

SESAME: 9:00 – 12:00 and 1:00 – 4:00 M-F Babai (Discrete): 4:15 – 6:00 TTh (SESAME participants: come on in) Bader (Geometry): 4:15 – 5:15 MWF Kerr (Algebraic geometry): 5:30 – 6:30 MWF May (For YSP participants): 11:00 – 12:00 MWF

EIGHTH WEEK:

SESAME: 9:00 – 12:00 and 1:00 – 4:00 M-F Babai (Discrete): 3:00 – 4:45 MWF (SESAME participants: come on in) Bader (Geometry): 4:30 – 6:00 TTh Kerr (Algebraic geometry): 5:00 – 6:00 MWF May (For YSP participants): 11:00 – 12:00 MWF

Tentative, August 11 and 12:

THURSDAY, 6:30 pm: Student presentations and dinner FRIDAY, 6:30 pm: Student presentations and dinner